OpenStudy (cutiecomittee123):

----cosb=1/2[cos(a+b)+cos(a-b)] cosa sinb sina cosb

OpenStudy (gorv):

can you expand cos(a+b) ???

OpenStudy (gorv):

or you know the cos C & D formula ???

OpenStudy (cutiecomittee123):

cos(a+b)=cosacosb-sinasinb

OpenStudy (gorv):

yes and cos(a-b)??

OpenStudy (cutiecomittee123):

cos(a-b)=cosacosb+sinasinb

OpenStudy (cutiecomittee123):

neither work because there isnt sin in the inital equation and it isnt adding up for me

OpenStudy (gorv):

now put these values in equation

OpenStudy (cutiecomittee123):

Im guessing its sina that is the answer just on estimation

OpenStudy (gorv):

replace cos(a+b) by its new value and similar with other ..

OpenStudy (gorv):

just follow you will learn dear...leave the ans it will come

OpenStudy (gorv):

can u replace ??

OpenStudy (cutiecomittee123):

whats its new value?

OpenStudy (gorv):

cos(a+b)=cosacosb-sinasinb

OpenStudy (gorv):

you know now what u need to use in place of cos(a+b)

OpenStudy (gorv):

got it ??

OpenStudy (gorv):

replace cos(a+b) by cosa cosb-sinasinb

OpenStudy (gorv):

\[\frac{ 1 }{ 2 }[cosa cosb -sina sinb +cosacosb+sinasinb]\]

OpenStudy (gorv):

yeah we just replace the identity by its expansion

OpenStudy (gorv):

so you gott it??

OpenStudy (cutiecomittee123):

yeah, but now what?

OpenStudy (gorv):

you see that sin terms can cancel each other??

OpenStudy (cutiecomittee123):

so 1/2[cosacosb+cosacosb]

OpenStudy (superdavesuper):

1/2[cosacosb+cosacosb] can be further simplified to....

OpenStudy (cutiecomittee123):

idk:( @superdavesuper

OpenStudy (superdavesuper):

cosacosb+cosacosb = 2cosacosb right?

OpenStudy (cutiecomittee123):

yes, sorry it took me so long to get back. Can you still walk me through this!??:) @superdavesuper

OpenStudy (superdavesuper):

@cutiecomittee123 plz re-read the posts above - its really all there :)

OpenStudy (cutiecomittee123):

so the answer is cosa